### v1

# and Machine Learning

Abstract
Denoising, the process of reducing random fluctuations in a signal to emphasize essential patterns, has been a fundamental problem of interest since the dawn of modern scientific inquiry. Recent denoising techniques, particularly in imaging, have achieved remarkable success, nearing theoretical limits by some measures. Yet, despite tens of thousands of research papers, the wide-ranging applications of denoising beyond noise removal have not been fully recognized. This is partly due to the vast and diverse literature, making a clear overview challenging.
This paper aims to address this gap. We present a comprehensive perspective on denoisers, their structure, and desired properties. We emphasize the increasing importance of denoising and showcase its evolution into an essential building block for complex tasks in imaging, inverse problems, and machine learning. Despite its long history, the community continues to uncover unexpected and groundbreaking uses for denoising, further solidifying its place as a cornerstone of scientific and engineering practice.
1 Introduction
Like all things of fundamental importance, image denoising is easy to describe, and very difficult to do well in practice. It is therefore not surprising that the field has been around since the beginning of the modern scientific and technological age - for as along as there have been sensors to record data, there has been noise to contend with.
Consider an image $\mathbf{x}$, composed of a “clean” (smooth [^1] It is important to note that this “smooth” component can contain edges and textures, hence we are using the term rather loosely here to describe operators that remove small-scale effects, leaving larger scale and higher contrast discontinuities alone.) component $\mathbf{u}$, and a “rough” or noisy component $\mathbf{e}$, which we take to be zero-mean Gaussian white noise of variance $\sigma^{2}$, going forward:
$\mathbf{x}=\mathbf{u}+\mathbf{e},$ (1)
where all images are scanned lexicographically into vectors. The aim of any denoiser is to decompose the image $\mathbf{x}$ back into its constituent components - specifically, to recover an estimate of $\mathbf{u}$, the underlying signal, by applying some operator (denoiser) $f(\cdot,\alpha)$, parameterized by some $\alpha$ as follows:
$\mathbf{\hat{x}(\alpha)}=f(\mathbf{x};\alpha)\approx\mathbf{u}.$ (2)
where $\alpha(\sigma^{2})$ is a monotonic function of the noise variance, and therefore controls the ”strength” of the denoiser.
As the description above indicates, a denoiser is not a single operator but a family of bounded [^2] We assume all images are in the numerical range [$0,1$]. In practice, an $8$-bit image would have values in [$0,255$] range. maps $f(\mathbf{x},\alpha):\left[0,1\right]^{N}\rightarrow\left[0,1\right]^{N}$. We expect ”good” denoisers to have certain naturally desirable properties, which alas in practice, many do not. For the sake of completeness, and as a later guide for how to design good denoisers, we call a denoiser ideal if it satisfies the following properties:
Property 1.
(Identity) When there is no noise (i.e. $\alpha=0$), the ideal denoiser will reproduce the input unchanged.
$f(\mathbf{x},0)=\mathbf{x},\;\;\;\;\;\;\forall\mathbf{x}$ (3)
That is, $f(\mathbf{x},0)$ is the identity operator.
Property 2.
(Conservation) An ideal denoiser has a symmetric Jacobian
$\nabla f(\mathbf{x},\alpha)=\nabla f(\mathbf{x},\alpha)^{T}$ (4)
Or equivalently,
$f(\mathbf{x},\alpha)=\nabla\mathcal{E}(\mathbf{x},\alpha)$ (5)
for some scalar-valued, differentiable (”potential” or ”energy”) function $\mathcal{E}(\mathbf{x},\alpha)$. This also means that the ideal denoiser defines a conservative vector field.
To convey some intuition for this property, consider the linear case. When a denoiser is linear: $f(\mathbf{x},\alpha)=W(\alpha)\mathbf{x}$, we always require the matrix $W(\alpha)$ to be row stochastic (meaning the rows sum to $1$) in order to preserve the mean local brightness. Ideally, we also require $W(\alpha)$ to be symmetric [1], which has the added advantage that the denoiser is admissible [2] in the mean-square sense. Property $2$ extends these notions to more general nonlinear denoisers  [^3] It’s worth noting that the combination of symmetric and row-stochastic implies that $W(\alpha)$ is doubly-stochastic..
Remark:
The conservation Property 2 guarantees that the ideal denoiser is the gradient of a scalar field. This also implies that $f(\mathbf{x},\alpha)$ is a Lipschitz map with some constant $M(\alpha)$:
$\|f(\mathbf{x},\alpha)-f(\mathbf{y},\alpha)\|\leq M(\alpha)\|\mathbf{x}-\mathbf{y}\|$ (6)
We naturally expect $f(\mathbf{0},\alpha)=\mathbf{0}$ for all $\alpha$; therefore, this Lipschitz condition is equivalent to $\|f(\mathbf{x},\alpha)\|\leq M(\alpha)\|\mathbf{x}\|$. A non-expansive denoiser would require that $M(\alpha)\leq 1$. In the statistics literature, such operators are called shrinking smoothers [3, 4].
The above properties impose the structure of an affine space [5] on the class of ideal denoisers. Namely, any affine combination of ideal denoisers is also ideal. That is, if we let [^4] Note that we do not place a constraint on the sign of $a_{k}$’s.
$g(\mathbf{x},\underline{\alpha})=\sum_{k=0}^{N}a_{k}\>f(\mathbf{x},\alpha_{k})\;\;\;\;\;\;\;\;\;\text{with}\;\;\;\;\;\sum_{k=0}^{N}\alpha_{k}=1$ (7)
it is easy to verify that Properties $1$ and $2$ are satisfied.
Now let’s consider the composition of ideal denoisers as follows:
$g_{n}(\mathbf{x},\underline{\alpha})=f_{1}(f_{2}(\cdots f_{n}(\mathbf{x},\alpha_{n}),\cdots,\alpha_{2}),\alpha_{1})$ (8)
Again, it is easily seen that Properties $1$ and $2$ are satisfied. We therefore have that the set of ideal denoisers is closed under composition.   Summary: Ideal denoisers satisfy:
  • Property 1:    $f(\mathbf{x},0)=\mathbf{x}$,
  • Property 2:    $f(\mathbf{x},\alpha)=\nabla\mathcal{E}(\mathbf{x},\alpha)$,
  • Closed-ness under affine linear combination.
  • Closed-ness under composition.
It is an unfortunate fact that in practice, most denoisers are not ideal. But this should not bother the reader, as by studying the broader class of denoisers we will learn how the above desirable properties are manifested or desired in practice, and which practical denoisers (approximately or exactly) satisfy them.
A note on this work: Rather than a survey of image denoising, this work focuses on defining ideal denoisers, their properties, and their connections to statistical theory and machine learning. We then demonstrate how these powerful components can serve as building blocks in various applications. Readers interested in a historical overview of image denoising are encouraged to consult the excellent resources in  [6, 7, 8, 9, 10].
2 Denoising as a Natural Decomposition
One of the remarkable aspects of well-behaved (even if not ideal) denoising operators is that we can employ them to easily produce a natural multi-scale decomposition of an image, with perfect reconstruction property [^5] This is similar in spirit to a multiscale decomposition, except that there is no decimation, and the filters are nonlinear here.. To start, consider a denoiser $f(\mathbf{x},\alpha)$. We can write the obvious relation:
$\mathbf{x}=f(\mathbf{x},\alpha)+\left[\mathbf{x}-f(\mathbf{x},\alpha)\right]$ (9)
The first term on the right-hand side is a smoothed (or denoised) version of $\mathbf{x}$, whereas the second term in the brackets is the residual $r_{0}(\mathbf{x},\alpha)=\mathbf{x}-f(\mathbf{x},\alpha)$ which is an ostensibly ”high-pass” version. Next, we can apply the same decomposition repeatedly to the already-denoised components [^6] To simplify the exposition, we use the same denoiser and the $\alpha$ at each step, but this is not necessary.:
$\mathbf{x}$ $=$ $f(f(\mathbf{x},\alpha),\alpha)+\left[f(\mathbf{x},\alpha)-f(f(\mathbf{x},\alpha),\alpha)\right]+r_{0}(\mathbf{x},\alpha)$ (10)
$=$ $f(f(\mathbf{x},\alpha),\alpha)+r_{1}(\mathbf{x},\alpha)+r_{0}(\mathbf{x},\alpha)$ (11)
$\vdots$ (12)
$=$ $f^{n}(\mathbf{x},\alpha)+\sum_{k=0}^{n-1}r_{k}(\mathbf{x},\alpha)$ (13)
where $f^{n}$ denotes the operator applied $n$ times (i.e. a diffusion process), and $r_{k}=f^{k}-f^{k+1}$ (i.e. a residual process). For any $n$, this $n$-th order decomposition splits $\mathbf{x}$ exactly into a smooth component $f^{n}(\mathbf{x},\alpha)$ and a sequence of increasingly fine-detail components $r_{k}(\mathbf{x},\alpha)$.
It is important to note that applying the operators $f(\mathbf{x},\alpha)$ multiple times does not necessary result in a completely smooth result. For instance, if we repeatedly apply a bilateral filter [11, 7], the result is a piece-wise constant image. The process we’ve described here has been called, in certain instances, a cartoon-plus-texture decomposition in [12, 13], mainly in the context of total-variation denoising. Our point of view is considerably more general, applicable to any denoiser.
Returning to the decomposition above, it empowers us to do practically useful things. For instance, truncating the residual terms at some $n$, we can smooth out certain high frequency features. More generally, we can null out any component in the sum; or better yet, recombine the components with new coefficients to produce a processed or modified image, as follows:
$g_{a}(\mathbf{x},\alpha)=\beta_{n}f^{n}(\mathbf{x},\alpha)+\sum_{k=0}^{n-1}\beta_{k}r_{k}(\mathbf{x},\alpha).$ (14)
An interesting side effect is that since ideal denoisers are closed under affine combination and composition, the resulting operators $g_{a}(\mathbf{x},\alpha)$ also inherit the ideal property. This approach was generalized and used in a practical setting in [14, 15] to produce a wide variety of image processing effects, built on a base of well-established (at the time) non-local means denoisers. This is illustrated in Figure 1. More recently, in [16] the authors used a similar decomposition to create a zero-shot method to control each individual component of the decomposition through diffusion model sampling.
Figure 1: Denoising as a natural image decomposition. Image adapted from[14].
Connection to Residual Networks:
The concept of breaking down an image into layers of varying detail is closely related to the architecture of Residual Neural Networks [17] (ResNets). Both share the principle that it’s simpler to model/learn residual mappings (the difference between the input and desired output) than to model/learn the complete transformation directly. While traditional deep neural networks try to learn this complex mapping in one go, ResNets use “skip connections” that allow the network to bypass layers, adding the original input to a later layer’s output. Let $H(\mathbf{x})$ be the desired complex mapping and $\mathbf{x}$ is the input, a ResNet layer attempts to learn a residual function $F(\mathbf{x})$ such that:
$H(\mathbf{x})=\mathbf{x}+F(\mathbf{x}),$ (15)
The skip connection ensures that the original input $\mathbf{x}$ is preserved and added back to the output. Note the connection to (9), where the residual term is exactly $-F(\mathbf{x})$. This decomposition and the use of skip connections simplify the network’s task, making optimization easier and mitigating the vanishing gradient problem that can hinder deep network training. Additionally, the preservation of the original input or its smooth approximation through skip connections ensures important information isn’t lost as data travels through the network [^7] There are actually ways of ensuring invertibility of ResNets, see e.g.,[18]. ResNets have been a major breakthrough in deep learning, enabling the training of much deeper networks and achieving state-of-the-art performance on image recognition tasks, with the concept of residual learning now being applied to other domains beyond image processing.
Image denoisers for anomaly detection:
The natural decomposition of an image using denoisers has also been used for analyzing images, for example to detect anomalies [19, 20] in images. The principle behind this is that anomalies, being infrequent occurrences, lack the self-similarity or smoothness typically observed in natural images. Drawing inspiration from patch-based denoising (e.g., non-local means), which employs self-similarity to differentiate between signal and noise, in [19] the authors introduce a method that effectively dissects an image into two components. The first is a self-similar component that embodies the background or ’smooth’ regions of the image given by the denoiser. The second is a residual component that encapsulates the distinctive, non-repetitive elements, which could potentially include anomalies and noise. The residual image, anticipated to resemble noise, is then subjected to a statistical test to detect any anomalies.
Next, we will describe various well-known classes of denoisers, including those derived from statistical optimality principles, and others which are pseudo-linear and derived from non-parametric or empirical considerations. We will also examine whether these classes of denoisers satisfy the above properties.
3 The Structure of General Denoisers
3.1 Bayesian Denoisers
Bayesian denoising invokes the use of a prior $P(\mathbf{u})$ on the class of “clean” images $\mathbf{u}$. This prior influences the estimate of the underlying signal away from the observed measurement $\mathbf{x}$. We will describe the popular Maximum a-Posteriori (MAP) and the Minimum Mean-Squared Error (MMSE) denoisers below.
The contrast between the MAP and MMSE is highlighted in Figure 2. The two estimates tend to coincide when the posterior is symmetric and unimodal, or when the noise variance $\sigma^{2}$ is small.
Figure 2: Bayesian Denoisers: MAP vs. MMSE.
Maximum a-Posteriori (MAP)
As the name indicates, the maximum a posteriori estimate is the value of $\mathbf{u}$ at which the posterior density $P(\mathbf{u}|\mathbf{x})$ is maximized,
[…]
When the noise is Gaussian and white, the optimization boils down to regularized least-squares
[…]
where $\phi(\mathbf{u})=-\log P(\mathbf{u})$ is the negative log-prior on the space of ”clean” images, and $\alpha$ is proportional to the noise variance. It would appear that the MAP denoiser does not have a closed form. However, the expression 17 is also known in the optimization literature as a proximal operator [21] when $\phi$ is convex, quasi-convex, or a difference of convex functions. It is well-known [22] that to every proximal operator $f$ there corresponds a convex scalar-valued function $\psi$ such that $f=\nabla\psi$.
Furthermore, in the context of the MAP estimate, $\psi$ has an explicit form:
[…]
where $\phi_{\alpha}$ is a smoothed version of $\phi$ called its Moreau envelope [21, 23]. As we will see below, the MMSE estimate shares a very similar form.
An example (for the scalar case) of the MAP denoiser for $\phi(\cdot)=\|\cdot\|_{1}$ is shown in Figure 3, where the resulting denoiser is exactly the soft-thresholding operator.
Figure 3: Example of MAP denoiser with $L_{1}$ loss, with $\alpha=1$. The Moreau envelope is the Huber loss.
Minimum Mean-Squared Error Denoising
In contrast to MAP denoising, MMSE denoising relies on the mean of the posterior density,
[…]
While the MMSE expectation integral is generally difficult or impossible to evaluate directly, a key result known as Tweedie’s formula [24, 25, 26] enables us to write the expression for MMSE also in the form of the gradient of a scalar function:
[…]
where $\tilde{\phi}_{\alpha}(\mathbf{x})=-\log P(\mathbf{x},\alpha)$. This is more or less identical to the form of the MAP denoiser in 18.
We graphically illustrate an example (for the scalar case) of the MMSE denoiser for $L_{1}$ loss in Fig. (5)(left). The two estimators are compared on the same plot in Fig 5 (right).
Figure 4: Illustration of a one-dimensional MMSE denoiser employing $L_{1}$ regularization, demonstrating the impact of varying $\alpha$. The visualization progresses from the smoothed distribution $P(\mathbf{x},\alpha)$ (left), to the corresponding Energy function (middle), and ultimately, the resulting denoiser (right).
Figure 5: Left: Example of MMSE denoiser with $L_{1}$ loss, with $\alpha=1$. Right: Comparison of MAP and MMSE denoiser for the $L1$ loss, with $\alpha=1$.
Data-driven MMSE denoisers: The typical modern supervised approach to image denoising is to train a deep neural network with pairs of clean and noisy images, where the noise is often modeled as additive white and Gaussian(AWGN) [CITE]. Let’s assume we have image pairs $(\mathbf{u},\mathbf{x})\sim P(\mathbf{u},\mathbf{x})$ where $\mathbf{u}$ represents a clean image, and $\mathbf{x}$ is the noisy observation obtained by adding AWGN with a known standard deviation to $\mathbf{u}$: $\mathbf{x}=\mathbf{u}+\mathbf{e}$, where $\mathbf{e}\sim\mathcal{N}(0,\sigma^{2})$.
A typical regression approach would predict $\mathbf{u}$ directly from $\mathbf{x}$ using a trained model $\hat{\mathbf{u}}(\mathbf{x})=F_{\theta}(\mathbf{x})\approx\mathbf{u}$, by minimizing the expected reconstruction error:
$\min_{\theta}\mathbb{E}_{\mathbf{u},\mathbf{e}}\|F_{\theta}(\mathbf{u}+\mathbf{e})-\mathbf{u}\|^{p}_{p}\approx\min_{\theta}\sum_{i}\|F_{\theta}(\mathbf{u}^{i}+\mathbf{e}^{i})-\mathbf{u}^{i}\|^{p}_{p}.$
[…]
As mentioned earlier, the MMSE denoiser is the average of all plausible clean signals given the noisy observation. This averaging can lead to a loss of details and a blurry appearance, especially when the noise level is high. Denoising Autoencoders (DAEs) are a prime example of data-driven MMSE denoisers [27]. These neural networks excel at learning robust data representations by training on noisy input and striving to reconstruct the original, clean data. This makes them not only valuable for denoising but also for tasks like data compression and feature extraction.
3.2 Energy-based Denoisers
We’ve seen that both MMSE and MAP estimators are of the form $f(\mathbf{x},\alpha)=\mathbf{x}-\alpha\nabla\phi_{\alpha}(\mathbf{x})$ where $\phi_{\alpha}(\mathbf{x})$ is some smoothed version of $\phi$ – differently smoothed in each case. These denoisers are also special cases of a general ”energy-based” formulation [28]:
$f(\mathbf{x},\alpha)=\nabla\mathcal{E}(\mathbf{x},\alpha)$ (22)
where in the particular case of MMSE and MAP,
$\mathcal{E}(\mathbf{x},\alpha)=\frac{1}{2}\|\mathbf{x}\|^{2}-\alpha\phi_{\alpha}(\mathbf{x})$ (23)
If the energy function satisfies $\nabla\mathcal{E}(\mathbf{x},0)=\mathbf{x}$ for all $\mathbf{x}$ (as do both the MMSE and MAP), then such denoisers are ideal. This is because the Jacobian of the denoiser can be written as
$\nabla f(\mathbf{x},\alpha)=\mathcal{H}\left[\mathcal{E}(\mathbf{x},\alpha)\right]$ (24)
where $\mathcal{H}$ denotes the Hessian operator which is, by definition, symmetric. In summary, all energy-based denoisers, including MAP and MMSE, are ideal, have symmetric Jacobians and are therefore conservative vector fields.
3.2.1 Approximation of Energy-based Denoisers
The energy-based formulation of denoisers provides a natural mechanism for principled empirical design of denoisers. This approach turns out to be consistent with the well-established empirical Bayes approach as well.
[…]
Now let’s write this marginal density in Gibbs form:
$P(\mathbf{x},\alpha)=\frac{1}{Z}\exp{\left[-\mathcal{E}(\mathbf{x},\alpha)\right]}.$ (25)
where $\mathcal{E}(\mathbf{x},\alpha)$ is an energy function with $\mathcal{E}(\mathbf{0},\alpha)=0$. The score function is related to the energy function as follows:
$s(\mathbf{x},\alpha)=\nabla_{\mathbf{x}}\log P(\mathbf{x},\alpha)=-\nabla_{\mathbf{x}}\mathcal{E}(\mathbf{x},\alpha).$ (26)
The smoothness of $P(\mathbf{x},\alpha)$ implies smoothness of the energy $\mathcal{E}(\mathbf{x},\alpha)$, thereby ensuring the existence of the gradient for both.
We can expand the energy around $\mathbf{x}=0$ using a first order Taylor expansion (with the Lagrange form of the remainder) to get
$\mathcal{E}(\mathbf{x},\alpha)$ $=\mathcal{E}(0,\alpha)+\nabla\mathcal{E}(0,\alpha)\mathbf{x}+\frac{1}{2}\mathbf{x}^{T}\mathbf{L}(\mathbf{a},\alpha)\mathbf{x}$ (27)
$=\nabla\mathcal{E}(0,\alpha)\mathbf{x}+\frac{1}{2}\mathbf{x}^{T}\mathbf{L}(\mathbf{a},\alpha)\mathbf{x},$
where $\mathbf{L}(\textbf{a},\alpha)$ represents the (symmetric) Hessian matrix of $\mathcal{E}$ evaluated at some (unknown) point $\mathbf{a}$ lying on the line segment [^8] As such, the point $\mathbf{a}$ depends indirectly on $\mathbf{x}$. between 00 and $\mathbf{x}$. Accordingly, the score function is
$s(\mathbf{x},\alpha)=-\nabla_{\mathbf{x}}\mathcal{E}(\mathbf{x},\alpha)=-\nabla\mathcal{E}(0,\alpha)-\mathbf{L}(\mathbf{a},\alpha)\mathbf{x}.$ (28)
Tweedie’s formula implies that the MMSE denoiser has the form:
$f(\mathbf{x},\alpha)$ $=\mathbf{x}-\alpha\>s(\mathbf{x},\alpha)$ (29)
$=\mathbf{x}+\alpha\nabla\mathcal{E}(0,\alpha)+\alpha\mathbf{L}(\mathbf{a},\alpha)\mathbf{x}.$
Requiring that $f(0,\alpha)=0$ implies that the second term must be zero. Therefore, the MMSE denoiser has a simple (pseudo [^9] Denoisers of this form are pseudo-linear [29] as they are similar in form to linear filters, except that the matrix $\mathbf{L}$ implicitly depends on $\mathbf{x}$.)-linear form:
[…]
To summarize, the resulting locally optimal denoiser can be written as
$f(\mathbf{x},\alpha)=\mathbf{W}(\mathbf{x},\alpha)\>\mathbf{x}$ (31)
where the symmetric matrix $\mathbf{W}(\mathbf{x},\alpha)$ is adapted to the structure of the input $\mathbf{x}$. This observation is consistent with earlier findings [7, 30, 31] that such pseudo-linear filters -including those built from (bias-free) deep neural nets- are (a) attempts at empirical approximations of the optimal MMSE denoiser, (b) shrinkage operations in an orthonormal basis adapted to the underlying structure of the image, and (c) perturbations of identity. In particular, such denoisers can be written in the form $f(\mathbf{x})=\nabla f(\mathbf{x})\mathbf{x}$, meaning that their local behavior is fully determined by their Jacobian, and therefore its spectrum.
Though these facts were neither historically clarified, nor the original motivation for their development [29], denoisers of the form 31)*** have always been heuristic/empirical approximations to the MMSE. These denoisers were hugely popular and effective (e.g. [32, 11]) for decades before the more recent introduction of neural networks.
Next, we will describe these types of denoisers -using the language of kernels- in more detail.
3.3 Kernel Denoisers
Motivation:
The basic idea behind kernel denoisers follows a non-parametric approach to modeling the distribution of (clean) images. Concretely, consider our basic setting
$\mathbf{x}=\mathbf{u}+\mathbf{e},$ (32)
where $\mathbf{e}$ is zero-mean Gaussian white noise of variance $\alpha=\sigma^{2}$. In practice the density $P(\mathbf{u})$ is unknown, but we may have access to examples [^10] Without loss of generality, $\mathbf{u}_{i}$ can refer to either full images, or patches thereof. $\mathbf{u}_{i}$, for $i=1,\cdots,n$. We can construct a naive empirical estimate of the distribution as follows:
$\widehat{P}(\mathbf{u})=\frac{1}{n}\sum_{i=1}^{n}\delta(\mathbf{u}-\mathbf{u}_{i})$
The empirical density for $\mathbf{x}$ is the convolution of $\widehat{P}(\mathbf{u})$ with the Gaussian density $\mathcal{N}(0,\alpha\mathbf{I})$, yielding:
$\widehat{P}(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})$
Armed with this estimate, we can compute an empirical estimate of the score:
$\nabla\log\widehat{P}(\mathbf{x})=\frac{-1}{\sigma^{2}}\left[\mathbf{x}-\frac{\sum_{i}\mathbf{u}_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}{\sum_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}\right].$ (33)
Invoking Tweedie’s formula, we have a closed form approximation to the MMSE denoiser as a (data-dependent) weighted average of the clean data points $\mathbf{u}_{i}$:
$\widehat{\mathbf{x}}_{mmse}$ $\approx\mathbf{x}+\sigma^{2}\nabla\log\widehat{P}(\mathbf{x})$ (34)
$=\mathbf{x}-\left[\mathbf{x}-\frac{\sum_{i}\mathbf{u}_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}{\sum_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}\right]$
$=\frac{\sum_{i}\mathbf{u}_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}{\sum_{i}\mathcal{N}(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})}$
$=\sum_{i}\mathbf{u}_{i}W(\mathbf{x}-\mathbf{u}_{i},\alpha\mathbf{I})$
In practice, we may only have access to the noisy image $\mathbf{x}$. In this scenario, we can treat each pixel $x_{i}$ as an independent sample (with independent noise) and apply the same reasoning directly to the noisy input, using it as a proxy for the clean signals:
$\hat{x}_{i}=\sum_{i}x_{j}W(x_{j}-x_{i},\alpha).$
This is a primitive instance of the pseudo-linear form alluded to earlier. In particular, the Gaussian ”kernels”, motivated by the assumed (Gaussian) distribution of the noise, can be thought of more generally as one of a myriad of choices of positive-definite kernels that can be employed to construct more general denoisers, as described below.
The General Pseudo-Linear Form
The pseudo-linear form is very convenient for the analysis of practical denoisers in general [7, 33]. But even more importantly, it is a fundamental and widespread approach to denoising that decomposes the operation into two distinct steps. First is a nonlinear step where data-dependent weights $\mathbf{W}(\mathbf{x},\alpha)$ are computed. Next is a linear step where weighted averages of the input pixels yield each output pixel. More specifically, for each output pixel $x_{i}$, the denoiser can be described as:
$\hat{x}_{i}=\sum_{j}W_{ij}(\mathbf{x},\alpha)\>x_{j}.$ (35)
Gathering all the weights into a matrix $\mathbf{W}(\mathbf{x},\alpha)$ reveals the denoiser in pseudo-linear matrix form:
$f(\mathbf{x},\alpha)=\mathbf{W(x,\alpha)}\>\mathbf{x}.$ (36)
Generally speaking, the weights are computed based on the affinity (or similarity) of pixels, measured using a ”kernel” (a symmetric positive-definite function). When properly normalized, these yield the weights used to compute the output pixels as a weighted average. For instance, in the non-local means [34] case
$K_{ij}(\mathbf{x},\alpha)=\exp(-\|\mathbf{R}_{ij}\mathbf{x}\|^{2}/2\alpha^{2}),\;\;\;\;\;\text{where}\;\;\;\;\|\mathbf{R}_{ij}\mathbf{x}\|^{2}=\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}$ (37)
and $\mathbf{x}_{i}$ denotes a patch of pixels centered at $i$. There exist many other possibilities [35], a practical few of which are shown in Table 1.
Table 1: Some well-known isotropic, positive-definite kernels $K(\|\mathbf{R}_{ij}\mathbf{x}\|,\alpha)$
Name Kernel
Gaussian $\exp(-\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}/\alpha)$
Exponential $\exp(-\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{1}/\alpha)$
Cauchy $1/\left(1+\alpha\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}\right)$
When normalized, these affinities give the weights $W_{ij}$ as follows
$W_{ij}=\frac{K_{ij}}{\sum_{j}K_{ij}}\;\;\;\;\;\;\implies\;\;\;\;\;\sum_{i}W_{ij}=1$ (38)
In more compact notation [^11] Since the weights sum to $1$ across the rows of $\mathbf{W}$, this matrix is row-stochastic. :
$\mathbf{W}(\mathbf{x},\alpha)=\mathbf{D}^{-1}(\mathbf{x},\alpha)\mathbf{K}(\mathbf{x},\alpha)$ (39)
where $\mathbf{D}(\mathbf{x},\alpha)=\operatorname{diag}[d_{1},d_{2},\cdots,d_{N}]$ is a diagonal normalization matrix constructed from the row sums ($d_{i}={\sum_{j}K_{ij}}$) of $\mathbf{K}(\mathbf{x},\alpha)$.
Remark:
For common kernels such as those highlighted in the above table, the parameter $\alpha$ controls the spread of the kernel. Therefore, as $\alpha\rightarrow 0$, the kernel approaches a scaled Dirac delta: $K_{ij}(\mathbf{x},0)=d\>\delta_{ij}$, or equivalently, the Kernel matrix is a scaled identity: $\mathbf{K}(\mathbf{x},0)=d\mathbf{I}$. Consequently, normalizing gives $\mathbf{W}(\mathbf{x},0)=\mathbf{I}$. If in addition $\mathbf{W}(\mathbf{x},\alpha)$ is symmetric, then the denoiser can be approximated as the gradient of an energy (see discussion in previous Section 3.2):
$f(\mathbf{x},\alpha)=\nabla\left[\mathbf{x}^{T}\mathbf{W}\mathbf{x}\right]$ (40)
In practice, symmetry of the filter matrix $\mathbf{W}(\mathbf{x},\alpha)$ is not a given [^12] Though one can empirically verify that such weight matrices are approximately symmetric [36, 37]. Despite the fact that the kernel matrix $\mathbf{K}(\mathbf{x},\alpha)$ is symmetric, the resulting weight matrix $\mathbf{W}(\mathbf{x},\alpha)=\mathbf{D}^{-1}(\mathbf{x},\alpha)\mathbf{K}(\mathbf{x},\alpha)$ is not so, due to the non-trivial diagonal normalization by $\mathbf{D}$. Fortunately, one can modify $\mathbf{W}(\mathbf{x},\alpha)$ to satisfying the symmetry condition as detailed in [1, 38]. This is accompliched by applying Sinkhorn balancing to $\mathbf{W}$ (or equivalently to $\mathbf{K}$), resulting in a symmetric and doubly-stochastic weight matrix, which can incidentally improve mean-squared error denoising performance over the baseline.
Alternatively, one can take a different approach via a first-order Taylor series [38, 39]:
$\mathbf{W}(\mathbf{x},\alpha)\approx\mathbf{I}+\beta\left(\mathbf{K}(\mathbf{x},\alpha)-\mathbf{D}(\mathbf{x},\alpha)\right),$ (41)
where $\beta^{-1}=\frac{1}{N}\sum_{i}d_{ii}$. The right-hand side is evidently symmetric.
3.4 Summary
The takeaway message from the above discussion is that denoisers we described share some important properties in common. Namely, they have the form $f(\mathbf{x})=\mathbf{x}-\alpha g(\mathbf{x})$ where $g$ is the gradient of some scalar function. Furthermore, they are:
Perturbation of the Identity: The ideal behavior of a denoiser when the noise is absent ($\alpha=0$) is to give the input image back, unchanged. This is what we identified as Propertry 1 in the introductory Section 1. We’ve seen that both Bayesian (MAP, and MMSE) denoisers, and their (ideal) empirical approximations satisfy this condition.
Shrinkage Estimators: The general form $f(\mathbf{x})=\mathbf{x}-\alpha v(\mathbf{x})$ can be interpreted as the ”trivial” denoiser $\mathbf{x}$ with a correction term $\alpha v(\mathbf{x})$ that pulls the components of the noisy input toward zero. It is remarkable that these denoisers have the same form as the original James-Stein estimator [40], where $\mathbf{x}$ was interpreted as the maximum-likelihood estimator, and $\alpha v(\mathbf{x})$ played the role of a Bayesian ”correction”. It has been observed [7, 30, 31] that such denoisers behave (at least locally) as shrinkage operations in an orthonormal basis adapted to the underlying structure of the image.
Gradient Descent on Energy: We noted that many denoisers can be written in the form $f(\mathbf{x})=\mathbf{x}-\alpha\nabla\mathcal{E}(\mathbf{x})$. It is obvious that the right-hand side defines one step in a steepest descent iteration. Repeated applications of a denoiser have the effect of marching toward the local stationary point of the energy.
Approximate Projection: It has been pointed out elsewhere [41] that if we accept the assertion that real-world images with $N$ pixels are approximately contained in low-dimensional manifolds of $\mathcal{R}^{N}$ [42], then adding noise is equivalent to orthogonal perturbation away from the manifold, and denoising is approximately a projection onto the manifold. In particular, for small noise, denoising is precisely a projection onto the local tangent of said manifold. As such, the work of denoising is essentially analogous to manifold learning.
4 Denoising, the Score Function, and Generative Modeling
A crucial link between denoising and the score function enables denoisers to enable learning complex probability distributions. In modeling real-world data, and images in particular, we are typically faced with a complex, high-dimensional probability density, $P(\cdot)$. Explicitly modeling such a distributions can be computationally intractable or extremely difficult. The score function, defined as the gradient of the log probability density, can provide a way through.
$s(\mathbf{x},\alpha)=\nabla_{\mathbf{x}}\log P(\mathbf{x},\alpha).$ (42)
Instead of modeling the distribution $P(\mathbf{x},\alpha)$ directly, we can learn, or approximate, the score function[43]. Denoising techniques are a way to implicitly learn the score function roughly as follows: an estimate of the score function around a “clean” image is obtained by corrupting it with noise, training a model to reconstruct the original clean image from the noisy version, and measuring the denoising residual:
$s(\mathbf{x},\alpha)\approx\frac{\mathbf{x}-f(\mathbf{x},\alpha)}{\alpha}.$ (43)
At first blush, it is not at all clear why this is a reasonable procedure. Yet there are a number of ways [44, 45, 43, 46, 47] to motivate this idea -perhaps none more direct than by using Tweedie’s formula introduced earlier in Eq. 20.
[…]
Rewriting this establishes a direct and exact relationship between score function and the MMSE denoiser:
[…]
Despite its elegance, the MMSE estimator is typically difficult to compute, or entirely inaccessible. Therefore as a proxy, often other denoisers are used, which may only be rough approximations of the MMSE (Eq. 43).
One can take a broader point of view by considering ideal denoisers:
$f(\mathbf{x},\alpha)=\nabla\mathcal{E}_{0}(\mathbf{x},\alpha).$
where $\mathcal{E}_{0}(\mathbf{x},\alpha)$ is of the form
$\mathcal{E}_{0}(\mathbf{x},\alpha)=\frac{1}{2}\|\mathbf{x}\|^{2}-\alpha\mathcal{E}(\mathbf{x},\alpha).$ (46)
Energy functions such as these can be learned [43, 48, 49, 28], and the resulting denoisers have the appealing form [^13] We remind the reader this includes the MMSE and MAP, but can be more general.:
$f(\mathbf{x},\alpha)=\mathbf{x}-\alpha\nabla\mathcal{E}(\mathbf{x},\alpha).$ (47)
Or equivalently
$\nabla\mathcal{E}(\mathbf{x},\alpha)=\frac{\mathbf{x}-f(\mathbf{x},\alpha)}{\alpha}.$ (48)
This illustrates again that the energy function is a proxy [48] for the score $s(\mathbf{x},\alpha)\approx\nabla\mathcal{E}(\mathbf{x},\alpha)$, and the resulting denoiser’s residual can be used as an approximation of the score.
4.1 Denoising as the Engine of Diffusion Models
Denoising Diffusion Models (DDMs) [45, 43, 50, 51, 47] have become an important area of research in generative modeling. They operate by progressively corrupting training data with noise until it’s indistinguishable from random noise, then learning to systematically reverse this corruption. By training a model to iteratively denoise, it gains the ability to generate entirely new, coherent data samples from a starting point of pure noise, effectively converting noise into meaningful structures like images or other data forms (Figure  6).
Figure 6: Diffusion: Forward (Modeling) and Backward (Sampling).
Despite their popularity, expressive power, and tremendous success in practice, there’s been relatively little intuitive clarity about how they operate. At their core, these models enable us to start with a sample from one distribution (e.g. a Gaussian), and arrive (at least approximately) at a sample from a target distribution $P(\mathbf{x},\alpha)$. But how is this magic possible? Referring to Figure 6, let’s say we begin with a sample $\mathbf{x}_{T}\sim\mathcal{N}(0,\alpha\mathbf{I})$, where $\alpha\gg\text{Var}[\mathbf{x}]$.
One simple way to activate this sampling process is to directly consider a flow differential equation
$\frac{d\mathbf{x}_{t}}{dt}=-\frac{1}{2}\frac{d\alpha_{t}}{dt}\nabla\log P(\mathbf{x}_{t},\alpha_{t}),$ (49)
where the right-hand side is the score function introduced earlier, and $\alpha_{t}$ is the noise level at time $t$. This differential equation, called a probability flow[47], by construction moves the initial condition gradually toward the distribution $P(\mathbf{x},\alpha)$. Solving this equation requires (a) selecting a numerical scheme, and (b) having access to the score function.
If we have access to an MMSE denoiser at every $t$, we can invoke Tweedie’s formula to write:
$\frac{d\mathbf{x}_{t}}{dt}$ $=\frac{1}{2}\frac{d\alpha_{t}}{dt}\frac{\left(\mathbf{x}_{t}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]\right)}{\alpha_{t}},$ (50)
which we call a residual flow. As we’ve described in the previous sections, lack of access to the MMSE denoiser forces us to select a different denoiser and therefore solve only an approximate version of the desired Eq. 49.
Understanding the Velocity Coefficient. A key question arises: How is the velocity coefficient (the term multiplying the residual) in (50) ODE determined? Let’s assume the process has a conditional variance $\text{Var}[\mathbf{x}_{t}|\mathbf{x}_{0}]=\alpha_{t}$ at time $t$ (i.e., noise level). The ODE is then constructed such that this variance evolves consistently, meaning $\text{Var}[\mathbf{x}_{t-dt}|\mathbf{x}_{0}]=\alpha_{t-dt}$.
Proof.
A first order discretization of (50) yields:
$\mathbf{x}_{t}-\mathbf{x}_{t-dt}=\frac{\alpha_{t}-\alpha_{t-dt}}{2\alpha_{t}}\left(\mathbf{x}_{t}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]\right).$ (51)
From this, we can derive:
$\text{Var}[\mathbf{x}_{t-dt}|\mathbf{x}_{0}]$ $=\frac{(\alpha_{t}+\alpha_{t-dt})^{2}}{4\alpha_{t}^{2}}\text{Var}[\mathbf{x}_{t}|\mathbf{x}_{0}]$ (52)
$=\frac{(\alpha_{t}+\alpha_{t-dt})^{2}}{4\alpha_{t}}$ (53)
$=\alpha_{t-dt}+\frac{(\alpha_{t-dt}-\alpha_{t})^{2}}{4\alpha_{t}}$ (54)
$\approx\alpha_{t-dt}.$ (55)
A crucial point for our discussion is that the probability flow described by equation (49) can be proven to yield the same marginal distributions as the stochastic formulation presented in [47]. This implies that, in the limit, if we initialize with samples from a Gaussian distribution, the solution is guaranteed to produce samples that match the data distribution. While a comprehensive mathematical analysis of diffusion models is beyond the scope of this work, we encourage interested readers to delve into the foundational works [45, 43, 50, 51, 47] or the excellent introductory overviews [52, 53, 54] for a deeper understanding.
5 Denoisers in the Context of Inverse Problems
Consider the following formulation of a linear inverse problem: The data is given by the following model $\mathbf{y}=H\mathbf{x}+\mathbf{e}$. A nominal solution can be obtained by solving this optimization problem:
$\hat{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{x}}l(\mathbf{y},\mathbf{x})+\lambda\mathbf{R}(\mathbf{x},\alpha)$ (56)
where $l(\mathbf{y},\mathbf{x})=\frac{1}{2}\|H\mathbf{x}-\mathbf{y}\|^{2}$ captures the Gaussian nature of the noise, and $\mathbf{R}(\cdot)$ is a regularization term intended to stabilize the solution, and $\lambda>0$ is a regularization parameter.
Over the last several decades, a vast number of choices for the regularizer $\mathbf{R}(\mathbf{x},\alpha)$ have been proposed with varying degrees of success. More recently, and independently of the machine learning literature, the development of strong denoising algorithms has had a profound impact on the field of inverse problems and low-level vision, as denoisers have been shown to enable powerful priors/regularizers [55, 36].
As we described earlier, access to a high quality denoising engine affords us the possibility to learn, or at least locally approximate, the geometry of the image manifold. This approximate geometry is learned based on a residual: the difference between a noisy image and its denoised version. This enables us to formulate inverse problems as general optimization tasks, where the denoiser (or more specifically a functional based on it) is used as a regularizer.
In order to solve the optimization problem (56), it is necessary to evaluate the gradient of the objective, which is as follows:
$H^{T}(\mathbf{y}-H\mathbf{x})+\lambda\nabla\mathbf{R}(\mathbf{x},\alpha)$ (57)
A key concern is how to compute $\nabla\mathbf{R}(\mathbf{x},\alpha)$. In this respect, classical choices of $\mathbf{R}(.)$ such as $L_{p}$ norms have been fairly convenient and successful; but also shown to have limited power to model natural images.
Another choice that has proved more effective is (image-adaptive) Laplacian regularizers [56, 57, 58, 59] that implicitly contain a (pseudo-linear) denoiser inside. Namely,
$\mathbf{R}(\mathbf{x},\alpha)=\frac{1}{2}\mathbf{x}^{T}\;\mathbf{L}(\mathbf{x},\alpha)\;\mathbf{x}$ (58)
In [36], we developed a natural extension of this idea called Regularization by Denoising (RED), where the regularizer is constructed from a more general denoiser:
$\mathbf{R}_{red}(\mathbf{x},\alpha)=\frac{1}{2}\mathbf{x}^{T}(\mathbf{x}-f(\mathbf{x},\alpha)).$ (59)
Note the intuition behind this prior: the value of $\mathbf{R}_{red}(\mathbf{x},\alpha)$ is low if the cross-correlation between the image and the denoising residual is small, or if the residual itself is small due to $\mathbf{x}$ being a fixed point of $f(\cdot)$.
But with this generality comes a challenge: can the gradient of the regularizer be computed easily? The answer is yes, when $f(\mathbf{x},\alpha)$ is ideal and locally homogeneous [36]. This is not difficult to prove:
$\nabla\mathbf{x}^{T}(\mathbf{x}-f(\mathbf{x},\alpha))$ $=$ $2\mathbf{x}-\nabla\left[\mathbf{x}^{T}f(\mathbf{x},\alpha)\right]$
$=$ $2\mathbf{x}-f(\mathbf{x},\alpha)-\nabla f(\mathbf{x},\alpha)\mathbf{x}$
$=$ $2\mathbf{x}-2f(\mathbf{x},\alpha)$
where the second line follows from the Jacobian symmetry of ideal denoisers; and the third line follows from local homogeneity and the definition of directional derivative [36]:
$\nabla f(\mathbf{x})\mathbf{x}=$ $\lim_{\epsilon\rightarrow 0}\frac{f(\mathbf{x}+\epsilon\mathbf{x})-f(\mathbf{x})}{\epsilon}$ (60)
$=$ $\lim_{\epsilon\rightarrow 0}\frac{(1+\epsilon)f(\mathbf{x})-f(\mathbf{x})}{\epsilon}=f(\mathbf{x})$ (61)
Replacing
$\nabla\mathbf{R}_{red}(\mathbf{x},\alpha)=\mathbf{x}-f(\mathbf{x},\alpha)$
for the gradient in 57, we have the following expression for the gradient of the objective:
$H^{T}(\mathbf{y}-H\mathbf{x})+\lambda(\mathbf{x}-f(\mathbf{x},\alpha))$
The most direct numerical procedure for solving this equation is a fixed point iteration that sets the gradient of the objective to zero:
$H^{T}(\mathbf{y}-H\mathbf{x}_{k+1})+\lambda(\mathbf{x}_{k}-f(\mathbf{x}_{k},\alpha))=0.$ (62)
Equivalently,
$\mathbf{x}_{k+1}$ $=\left[H^{T}H+\lambda I\right]^{-1}H^{T}\mathbf{y}+\lambda\left[H^{T}H+\lambda I\right]^{-1}f(\mathbf{x}_{k},\alpha)$ (63)
$\mathbf{x}_{k+1}$ $=\mathbf{b}+\mathbf{M}\;f(\mathbf{x}_{k},\alpha)$ (64)
where $\mathbf{b}$ is the (fixed) linear pseudo-inverse solution and $\mathbf{M}$ is also a fixed matrix. Procedurally, we start with $\mathbf{x}_{0}=\mathbf{y}$, denoise it, and then a linear operator $\mathbf{M}$ is applied and a bias is added - this leads to an updated estimate, and the process is repeated. Note that the structure of this iterative process is not altogether different from a denoising diffusion process[50] where a denoiser is repeatedly applied. In fact, when $H=I$ we see the structure of a bridge diffusion process [60]:
$\mathbf{x}_{k+1}$ $=\frac{1}{1+\lambda}\mathbf{y}+\frac{\lambda}{1+\lambda}f(\mathbf{x}_{k},\alpha)$ (65)
In a general statistical setting, the scalar valued $\mathbf{R}(\mathbf{x},\alpha)$ is often the result of assuming a prior whose negative-log is interpreted as the regularizer:
$\mathbf{R}(\mathbf{x},\alpha)=-\log P(\mathbf{x},\alpha),$ (66)
However, in cases where a denoiser is used to construct a regularizer, the role of the regularizer $\mathbf{R}(\mathbf{x},\alpha)\geq 0$ is that of an energy function that we implicitly use to define a Gibbs distribution [61]:
$P(\mathbf{x},\alpha)\propto\exp{\left[-\mathbf{R}(\mathbf{x},\alpha)\right]},$ (67)
In the particular case of RED: $\mathbf{R}_{red}(\mathbf{x},\alpha)=\frac{1}{2}\mathbf{x}^{T}(\mathbf{x}-f(\mathbf{x},\alpha))$, Equation 60 implies that an ideal and locally homogeneous denoiser has the form $f(\mathbf{x})=\nabla f(\mathbf{x})\mathbf{x}$, which means that under these conditions, the RED regularization can be thought of as a (pseudo-quadratic) energy function:
$\mathbf{R}_{red}(\mathbf{x},\alpha)=\frac{1}{2}\mathbf{x}^{T}\nabla f(\mathbf{x},\alpha)\mathbf{x}.$ (68)
Posterior Sampling with Denoisers
An alternative approach to solving inverse problems involves generating samples from the posterior distribution. In other words, given measurements $\mathbf{y}$, we aim to generate samples $\mathbf{x}$ that follow the distribution $P(\mathbf{x}|\mathbf{y})$. We can achieve posterior sampling from $P(\mathbf{x}|\mathbf{y})$ by modifying equation (49) as follows:
$\frac{d\mathbf{x}_{t}}{dt}=-\frac{1}{2}\frac{d\alpha_{t}}{dt}\nabla\log P(\mathbf{x}_{t}|\mathbf{y},\alpha_{t})=-\frac{1}{2}\frac{d\alpha_{t}}{dt}\left(\nabla\log P(\mathbf{x}_{t},\alpha_{t})+\nabla\log P(\mathbf{y}|\mathbf{x}_{t},\alpha_{t})\right),$ (69)
starting from $x_{T}\sim P(x_{T})$. The second equality is given by Bayes rule. We recognized the score function $\nabla\log P(\mathbf{x}_{t},\alpha_{t})$, that can be again related to the MMSE denoiser using Tweedie’s formula (20). The second term captures how much the current $\mathbf{x}_{t}$ explains the measurements $\mathbf{y}$ and it is generally intractable. As an example, let us briefly present DPS [62]. In [62], authors propose to approximate the intractable term with $\log P(\mathbf{y}|\mathbf{x}_{t})\simeq\log P(\mathbf{y}|\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}])$, which is a consequence of assuming $p(\mathbf{x}_{0}|\mathbf{x}_{t})\simeq\delta(\mathbf{x}_{0}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}])$. This leads to a tractable formulation:
$\frac{d\mathbf{x}_{t}}{dt}=-\frac{1}{2}\frac{d\alpha_{t}}{dt}\left(\frac{\mathbf{x}_{t}-\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]}{\alpha_{t}}+\rho_{t}\nabla\|H\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]-\mathbf{y}||^{2}\right),$ (70)
where $\rho_{t}$ is a hyperparameter controlling the balance between the prior and the measurements.
Denoising diffusion models are rapidly gaining attention as a powerful tool for solving inverse problems in various fields. This is often achieved by combining the strengths of diffusion models with additional approximations or techniques, as evidenced by a growing body of research (e.g., [63, 47, 64, 65, 66, 67, 68, 69, 70, 71]).
6 Conclusions
In this paper, we have explored the multifaceted nature of denoising, showcasing its far-reaching impact beyond the traditional task of noise removal. We have highlighted the structural properties of denoisers, their connection to Bayesian estimation and energy-based models, and their ability to act as powerful priors and regularizers in various applications. The surprising effectiveness of denoisers in tasks from generative modeling to inverse problems underscores their versatility and potential for future research. The continued evolution of denoising techniques, coupled with advancements in machine learning, promises to unlock even more innovative applications and deeper insights into the underlying structure of images.